Air (1100 K/ 101.32 kPa) is passing through a converging nozzle and leave it at 300 K. Determine the velocity of air at the nozzle outlet. The nozzle is laid horizontal. The inlet velocity of air can be ignored.


1
Expert's answer
2021-06-25T14:41:03-0400

"h_1+\frac{V_1^2}{2}=h_2+\frac{V_2^2}{2}"

But "V_1=0"

So, "V_2 = \sqrt{2(h_1-h_2)}"

"V_2= \sqrt{2c_p(T_1-T_2)}"

"V_2= \sqrt{2*1000(1100-300)}"

"V_2= 1265 m\/s"